### v1

@@ Abstract
For the numerical solution of nonsmooth problems, sometimes it is not necessary that an exact subgradient/generalized Jacobian is at our disposal, but that a certain semismoothness property is fulfilled. In this paper we consider not only semismoothness of nonsmooth real- and vector-valued mappings, but also its interplay with the semismoothness property for multifunctions. In particular, we are interested in the semismoothness of solution maps to parametric semismooth inclusions. Our results are expressed in terms of suitable generalized derivatives of the set-valued part, i.e., by limiting coderivatives or by SC (subspace containing) derivatives. As a byproduct we identify a class of multifunctions having the remarkable property that they are strictly proto-differentiable almost everywhere (with respect to some Hausdorff measure) on their graph.
@@ 1 Introduction
$\min\theta(x)\quad\mbox{ subject to }\quad x\in U_{\rm ad},$
where $\theta:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is a locally Lipschitzian function and $U_{\rm ad}\subset\mathbb{R}^{n}$ is a closed set, the use of bundle methods is a popular approach. There exists a bunch of bundle methods in the literature, cf. [17, 14, 29], and all these methods have in common that they require an
oracle, which returns for arbitrary $x\in\mathbb{R}^{n}$ the function value $\theta(x)$ and one subgradient $g(x)\in\partial_{c}\theta(x)$, (1)
where $\partial_{c}\theta$ denotes the Clarke subdifferential. Then, under the additional assumption that
$\theta$ is weakly semismooth,
it can be shown that the bundle method either stops after finitely many steps at a C-stationary point $\bar{x}$ or, if the produced sequence $x_{k}$ remains bounded, one of its accumulation point is C-stationary.
It seems to be that Dempe and Bard [4] were the first who observed that the Bundle-Trust region algorithm (BT algorithm) of Schramm and Zowe [29] also works, when the oracle only delivers a so-called pseudosubgradient. Pseudosubgradients were introduced by Norkin [22, 23] and have the remarkable feature that they fulfill a certain semismoothness property. Although this property does not necessarily imply that $\theta$ is weakly semismooth, one can show that a similar convergence result holds as above: If the BT-algorithm produces a bounded sequence then at least one accumulation point is pseudo-stationary, i.e., 00 is a pseudosubgradient of $\theta$ at this point. By checking the convergence proof for other bundle methods [14, 17] one can show that a similar assertion hold: If the oracle delivers only a pseudosubgradient, one can still prove that the method works and returns pseudo-stationary points. This is important for applications: Whereas the calculus for the Clarke subdifferential often yields only inclusions, the calculus for pseudosubdifferentials yields again a pseudosubdifferential.
@@
where $G:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is a locally Lipschitz mapping. Whereas in the original semismooth Newton method introduced by Qi and Sun [26] matrices from the Clarke generalized Jacobian were used, it was soon observed that the Clarke generalized Jacobian can be replaced by another mapping whose elements fulfill a certain semismoothness property closely related to the one imposed on pseudosubgradients. This is particularly important for carrying over the semismooth Newton method to the infinite-dimensional setting, where no generalized Jacobian exists. We refer the reader to [31] for the semismooth Newton method in the infinite-dimensional setting. Though the semismoothness property imposed on the pseudosubgradients and the matrices used in the semismooth Newton method is essentially the same, there is a big difference: Whereas for the semismooth Newton method the semismoothness needs to hold only at the solution, for the pseudosubdiiferentials it has to hold at every $x\in\mathbb{R}^{n}$, which has far-reaching consequences.
Finally, in [7] the semismooth Newton method for solving inclusions
$0\in F(x)$
was introduced, where $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ is a set-valued mapping. Here, the semismooth property was defined in terms of the limiting (Mordukhovich) coderivative and was later simplified to SC (subspace containing) derivative in [8]. The SC derivative can be considered as a generalization of the B-Jacobian (Bouligand) of single-valued mappings to multifunctions and acts as a kind of skeleton for the limiting coderivative. The semismooth Newton method has been already successfully applied to several difficult problems, see, e.g., [10, 11, 12].
We have seen that semismoothness plays an important role in solving different types of nonsmooth problems: optimization problems, equations and inclusions. The used notions of semismoothness differ slightly for the different problem types and so one aim of this paper is to introduce a unified notion which is compatible across the different problem types. This is important for the numerical solution of problems which are assembled by different nonsmooth objects. As an example consider an optimization problem of the form
@@
where $\varphi:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}$ is a locally Lipschitz function, $F:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{m}$ is a set-valued mapping and $\omega\subset\mathbb{R}^{n}$ is a closed set. Problems of such type arise, e.g., in bilevel games and mathematical programs with equilibrium constraints (MPECs).
@@
One possibility for tackling (2) is the so-called implicit programming approach (ImP), cf. [18]. Here we assume that for every $x$ the inclusion $0\in F(x,\cdot)$ has a unique solution denoted by $\sigma(x)$. Hence we can eliminate $y$ from the model above and arrive at the problem
$\min\,$ $\theta(x):=\varphi(x,\sigma(x))$
subject to $x\in\omega.$
@@
Under the additional assumption that $\sigma$ is locally Lipschitz, the objective $\theta$ is locally Lipschitz as well and we can apply a bundle method for its numerical solution, provided a suitable oracle is available. We cannot guarantee that we actually compute a Clarke subgradient of $\theta$; we will show, however, that we can determine a pseudosubgradient, provided the original data $\varphi$ and $F$ are semismooth and semismooth, respectively.
In this paper we will provide the theoretical base for this approach: In particular, we will give all the necessary definitions and provide the basic calculus rules. We prove the semismooth property of the solution maps of inclusions as appearing in (2) by the semismooth property of the underlying mapping $F$. Up to now, only results on the semismooth property of single-valued solution mappings implicitely given by single-valued equations are available, cf. [30, 24, 13, 19, 16]
Moreover, we will also shed light on a remarkable property of graphically Lipschitzian semismooth set-valued mappings, namely that they are strictly proto-differentiable at almost all points of its graph. This is in accordance with the fact that semismooth single-valued maps are almost everywhere strictly differentiable.
@@
Let us mention that for single-valued mappings there is a close relation between the semismoothness property and conservative derivatives, the latter being explored in a recent paper by Bolte and Pauwels [1]. In fact, it was shown in [6] that both notions are equivalent for semialgebraic maps. As semismooth derivatives also obey a chain rule like conservative Jacobians, the considerations about automatic differentiation from [1] carry over to our setting. Recently, a conservative Jacobian for solution maps given implicitely by equations was established in [2]. In [3], a conservative Jacobian for the solution map of a parametric monotone inclusion was obtained by rewriting the inclusion as an equivalent fixed-point equation for the forward-backward map by means of the resolvent. Though the resolvents are explicitly known for a lot of monotone mappings, it is not clear how a conservative Jacobian of the resolvent can be computed in the general case. In fact, the resolvent itself may be considered as the solution map of a shifted inclusion. On the contrary, our approach is based on the differentiation of the underlying set-valued mapping by SC derivatives.
Whereas this paper is of theoretical nature, in a subsequent paper we will demonstrate how we can apply the theory to several instances of problem (2) and will also provide some numerical experiments.
@@
The paper is organized in the following way. Section 2 is subdivided into several subsections, the first two containing the basic notions of variational analysis and generalized differentiation. In Subsection 2.3 we describe a general principle, analogous to the construction of Clarke generalized Jacobian, how to generate from an arbitrary multifunction another multifunction having some favorable properties. In the next subsection we recall the basic properties of SCD (subspace containing derivative) mappings and give in Theorem 2.18 a characterization of strict proto-differentiability. In Subsection 2.5 we recall the semismooth property of set-valued mappings and its extension to SCD mappings.
@@
Section 3 is devoted to the semismoothness property of single-valued mappings with respect to certain multifunctions. This concept is commonly used in semismooth Newton methods for solving nonsmooth equations, in particular in infinite dimensions. However, in contrast to these applications we do not consider the case that semismoothness only holds at the solution, but on the whole domain. There is a close relation with the seminal work of Norkin [22, 23] on generalized differentiation of scalar functions. In the next section we present some relations between the semismooth and the semismooth property for a class of multifunctions, whose graphs coincide locally after a change of coordinates with the one of a Lipschitzian mapping. Section 5 concerns the semismooth property of solution maps to inclusions. We establish the semismooth property of general solution maps and the semismooth property for continuous localizations and selections, respectively. We further demonstrate that the analysis considerably simplifies for SCD mappings.
@@
The following notation is employed. Given a linear subspace $L\subseteq\mathbb{R}^{n}$, $L^{\perp}$ denotes its orthogonal complement. Further, given a multifunction $F$, $\mathrm{gph}\,F:=\{(x,y)\,\big{|}\,y\in F(x)\}$ stands for its graph. For an element $u\in\mathbb{R}^{n}$, $\|u\|$ denotes its Euclidean norm and ${\cal B}_{\delta}(u)$ denotes the closed ball around $u$ with radius $\delta$. The closed unit ball is denoted by ${\cal B}$. In a product space we use the norm $\|(u,v)\|:=\sqrt{\|u\|^{2}+\|v\|^{2}}$. The space of all $m\times n$ matrices is denoted by $\mathbb{R}^{m\times n}$. For a matrix $A\in\mathbb{R}^{m\times n}$ , we employ the operator norm $\|A\|$ with respect to the Euclidean norm and we denote the range of A by ${\rm rge\;}A$. Given a set $\Omega\subset\mathbb{R}^{s}$, we define the distance of a point $x$ to $\Omega$ by ${\rm dist}(x,\Omega):=\inf\{\|y-x\|\,\big{|}\,y\in\Omega\}$ and the indicator function is denoted by $\delta_{\Omega}$. Further we define for a mapping $F:\Omega\to\mathbb{R}^{m}$, $\Omega\subset\mathbb{R}^{n}$ open, and $x\in\Omega$ the moduli
@@ 2.1 Nonsmooth analysis
Let us recall the notions of upper (outer) and lower (inner) limits in the sense of Painlevé–Kuratowski of a parameterized family $C_{t}$ of subsets of a metric space, where $t$ also belongs to a metric space:
$\mathop{{\rm Lim}\,{\rm sup}}_{t\to\bar{t}}C_{t}:=\{x\,\big{|}\,\liminf_{t\to\bar{t}}\,{\rm dist}(x,C_{t})=0\},$
$\mathop{{\rm Lim}\,{\rm inf}}_{t\to\bar{t}}C_{t}:=\{x\,\big{|}\,\limsup_{t\to\bar{t}}\,{\rm dist}(x,C_{t})=0\}.$
It easily follows from the definition that both the upper and lower limit sets are closed. These sets can also be described in terms of sequences as follows. The upper limit can be defined as the set of points $x$ for which there exists a sequence $t_{k}\to\bar{t}$ such that $x_{k}\to x$ for some $x_{k}\in C_{t_{k}}$ $\forall k$. Similarly, the lower limit can be defined as a set of points $x$ such that for every sequence $t_{k}\to\bar{t}$ we can find $x_{k}\in C_{t_{k}}$ such that $x_{k}\to x$. Clearly, $\mathop{{\rm Lim}\,{\rm inf}}_{t\to\bar{t}}C_{t}\subset\mathop{{\rm Lim}\,{\rm sup}}_{t\to\bar{t}}C_{t}$.
@@
If $F$ is locally Lipschitz, by Rademacher’s Theorem the set $\Omega\setminus{\cal O}_{F}$ is negligible, i.e., has Lebesgue measure 00. Since the Jacobians $\nabla F(x)$, $x\in{\cal O}_{F}$, are also locally bounded by the Lipschitz constant, we obtain ¯$\overline{\nabla}F(x)\subset\mathrm{conv}\,\overline{\nabla}F(x)$, $x\in\Omega$.
We say that $F$ is strictly differentiable at $\bar{x}\in\Omega$ if there is an $m\times n$ matrix $A$ such that
$\lim_{{\textstyle{x,x^{\prime}\to\bar{x}\atop x\not=x^{\prime}}}}\frac{F(x^{\prime})-F(x)-A(x^{\prime}-x)}{\|x^{\prime}-x\|}=0.$
Of course, this implies that $F$ is differentiable at $\bar{x}$ and $\nabla F(\bar{x})=A$. Strict differentiability of $F$ at $\bar{x}$ is equivalent with the requirements that $F$ is Lipschitz near $\bar{x}$ and
$\lim_{{\textstyle{x\to\bar{x}\atop t\downarrow 0}}}\frac{F(x+tu)-F(x)}{t}=Au,\ u\in\mathbb{R}^{n}$
for some $m\times n$ matrix $A$.
For a locally Lipschitzian function $f:\Omega\to\mathbb{R}$, $\Omega\subset\mathbb{R}^{n}$ open, we can define the Clarke subdifferential as
$\partial_{c}f(x):=\mathrm{conv}\,\overline{\nabla}f(x)^{T}:=\{g^{T}\,\big{|}\,g\in\mathrm{conv}\,\overline{\nabla}f(x)\},\ x\in\Omega.$
The elements of $\partial_{c}f(x)$ are called subgradients and are column vectors.
Given any extended real-valued function $f:\mathbb{R}^{n}\to(-\infty,\infty]$, we always assume that $f$ is proper, i.e., $f(x)>-\infty$ for all $x$ and $\mathrm{dom}\,f:=\{x\in\mathbb{R}^{n}\,\big{|}\,f(x)<\infty\}\not=\emptyset$. Given a point $\bar{x}\in\mathrm{dom}\,f$, the regular (Fréchet) subdifferential of $f$ at $\bar{x}$ is given by
^$\widehat{\partial}f(\bar{x}):=\Big{\{}x^{*}\in\mathbb{R}^{n}\,\big{|}\,\liminf_{x\to\bar{x}}\frac{f(x)-f(\bar{x})-\langle x^{*},x-\bar{x}\rangle}{\|x-\bar{x}\|}\geq 0\Big{\}},$
while the (limiting) subdifferential is defined by
$\partial f(\bar{x}):=\{x^{*}\,\big{|}\,\exists(x_{k},x_{k}^{*}){\mathop{\longrightarrow}\limits^{\mathrm{gph}\,\widehat{\partial}f}}(\bar{x},x^{*})\mbox{ with }f(x_{k})\to f(\bar{x})\}.$
If $\mathrm{dom}\,f=:\Omega$ is open, then clearly ${\cal O}_{f}\subset\mathrm{dom}\,\widehat{\partial}f$ and ¯$\overline{\nabla}f(\bar{x})^{T}\subset\partial f(\bar{x})$, provided that $f$ is continuous at $\bar{x}\in\Omega$. If $f$ is locally Lipschitz on $\Omega$ then
$\partial_{c}f(x)=\mathrm{conv}\,\partial f(x),\ x\in\Omega,$
cf. [28, Theorem 9.61].
@@
    $\lim\limits_{x{\mathop{\longrightarrow}\limits^{{\cal O}_{F}}}\bar{x}}\nabla F(x)$ exists.
@@
ad (i) $\Rightarrow$ (iii): If $F$ is strictly differentiable at $\bar{x}$ then $\nabla F$ is continuous at $\bar{x}$ relative to ${\cal O}_{F}$, cf. [28, Exercise 9.64]. Thus $\lim\limits_{x{\mathop{\longrightarrow}\limits^{{\cal O}_{F}}}\bar{x}}\nabla F(x)$ exists (and equals to $\nabla F(\bar{x})$).
@@
ad (ii) $\Rightarrow$ (i): Let ¯$\overline{\nabla}F(\bar{x})=\{A\}$. For every $i=1,\ldots,m$, we conclude from [5, Theorem 2.6.6] that the $i$-th component of $F$, $i=1,\ldots m$, satisfies ¯$\bar{\nabla}F_{i}(\bar{x})=\{A_{i}\}$, where $A_{i}$ denotes the $i$-th row of $A$. Thus $\partial_{c}F_{i}(\bar{x})$ is a singleton and therefore $F_{i}$ is strictly differentiable at $\bar{x}$ by [5, Proposition 2.2.4]. Since every component $F_{i}$, $i=1,\ldots m$, is strictly differentiable at $\bar{x}$, so is $F$ as well. ∎

### v2

@@ Abstract
For the numerical solution of nonsmooth problems, sometimes it is not necessary that an exact subgradient/generalized Jacobian is at our disposal, but it suffices that a semismooth derivative, i.e., a mapping satisfying a certain semismoothness property, is available. In this paper we consider not only semismooth derivatives of single-valued mappings, but also its interplay with the semismoothness property for multifunctions. In particular, we are interested in semismooth derivatives of solution maps to parametric semismooth inclusions. Our results are expressed in terms of suitable generalized derivatives of the set-valued part, i.e., by limiting coderivatives or by SC (subspace containing) derivatives. Further we show that semismooth derivatives coincide a.e. with generalized Jacobians and state some consequences concerning strict proto-differentiability for semismooth multifunctions.
@@ 1 Introduction
$\min\theta(x)\quad\mbox{ subject to }\quad x\in U_{\rm ad},$ (1.1)
where $\theta:\mathbb{R}^{n}\to\mathbb{R}$ is a locally Lipschitz function and $U_{\rm ad}\subset\mathbb{R}^{n}$ is a closed set, the use of bundle methods is a popular approach. There exists a bunch of bundle methods in the literature, cf. [23, 19, 36], and all these methods have in common that they require an
oracle, which returns for arbitrary $x\in\mathbb{R}^{n}$ the function value $\theta(x)$ and one subgradient $g(x)\in\partial^{c}\theta(x)$, (1.2)
where $\partial^{c}\theta$ denotes the Clarke subdifferential. Then, under the additional assumption that
$\theta$ is weakly semismooth (see Remark 3.10 below),
it can be shown that the bundle method either stops after finitely many steps at a C-stationary point $\bar{x}$ or, if the produced sequence $x_{k}$ remains bounded, one of its accumulation points is C-stationary.
From a limited number of bundle methods, applicable to the numerical solution of (1.1) (e.g. [23, 19, 36]), we consider in this paper the Bundle-Trust region algorithm (BT algorithm) from Schramm and Zowe [36], described also in [40]. It seems that Dempe and Bard [8] were the first who observed that the BT algorithm also works, when the oracle delivers only a so-called pseudogradient. Pseudogradients were introduced by Norkin [29, 30] and have the remarkable feature that they fulfill a certain semismoothness property. Although this property does not necessarily imply that $\theta$ is weakly semismooth, one can show that a similar convergence result holds as above: If the BT-algorithm produces a bounded sequence then at least one accumulation point is pseudo-stationary, i.e., 00 is a pseudogradient of $\theta$ at this point. By checking the convergence proof for other bundle methods [19, 23] one can show that a similar assertion holds: If the oracle delivers only a pseudogradient, one can still prove that the method works and returns pseudo-stationary points.
@@
where $G:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is a locally Lipschitz mapping. Whereas in the original semismooth Newton method, introduced by Kummer [22] and Qi and Sun [34], matrices from the Clarke generalized Jacobian were used, it was soon observed that the Clarke generalized Jacobian can be replaced by another mapping which fulfills a certain semismoothness property, closely related to the one imposed on pseudogradients. In what follows the elements of this mapping, both in the scalar- and vector-valued case, will be termed semismooth derivatives. This is particularly important for carrying over the semismooth Newton method to the infinite-dimensional setting, where no generalized Jacobian exists. We refer the reader to [38] for the semismooth Newton method in the infinite-dimensional setting. Note that, whereas the Clarke calculus often yields only inclusions, the calculus of semismooth derivatives yields again a semismooth derivative. Though the semismoothness property, imposed on the semismooth derivatives in the scalar- and vector-valued case, is essentially the same, there is a big difference: Whereas for the semismooth Newton method the semismoothness needs to hold only at the solution, when working with pseudogradients in the framework of a bundle method it needs to hold at each $x\in U_{ad}$.
Finally, for the numerical solution of inclusions
$0\in F(x),$
the so-called semismooth Newton method was introduced in [11], where $F:\mathbb{R}^{n}\rightrightarrows\mathbb{R}^{n}$ is a set-valued mapping. Here, the semismooth property was defined in terms of the limiting (Mordukhovich) coderivative and was later simplified to SC (subspace containing) derivative in [12]. The SC derivative can be considered as a generalization of the B-Jacobian (Bouligand) of single-valued mappings to multifunctions and acts as a kind of skeleton for the limiting coderivative. The semismooth Newton method has been already successfully applied to several difficult problems, see, e.g., [14, 15, 9, 10].
We observe that semismoothness plays an important role in solving different types of nonsmooth problems: optimization problems, equations and inclusions. The respective notions of semismoothness are slightly different for the different problem types, which complicates the numerical solution of problems assembled by different nonsmooth and set-valued objects. As an example consider an optimization problem of the form
@@
where $\varphi:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}$ is a locally Lipschitz function, $F:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightrightarrows\mathbb{R}^{m}$ is a set-valued mapping and $U_{\rm ad}\subset\mathbb{R}^{n}$ is a closed set. Problems of such type arise, e.g., in bilevel programming and mathematical programs with equilibrium constraints (MPECs).
@@
One possibility for tackling (1.3) is the so-called implicit programming approach (ImP), cf. [24], where one typically requires that for every $x$ the inclusion $0\in F(x,\cdot)$ has a unique solution denoted by $\sigma(x)$. Hence we can eliminate $y$ from the model above and arrive at the (reduced) problem
$\min\theta(x):=\varphi(x,\sigma(x))\quad\mbox{ subject to }x\in U_{\rm ad}.$
@@
Under the additional assumption that $\sigma$ is locally Lipschitz, the objective $\theta$ is locally Lipschitz as well and we can apply a bundle method for its numerical solution, provided a suitable oracle is available. In most cases, however, one cannot guarantee that at all points a Clarke subgradient of $\theta$ can be computed. Nevertheless, if the original data $\varphi$ and $F$ are semismooth and semismooth, respectively, then it is possible to construct an oracle returning an element of a semismooth derivative of $\theta$. This oracle will then ensure the convergence of the used bundle method to a stationary point.
In this paper we will provide the theoretical base for this approach. In particular, we will conduct a deep analysis of various available notions of the semismoothness property and find out some important mutual relationships. We will show that a single-valued mapping possesses some semismooth derivative on an open set $\Omega$ if and only if it is G-semismooth (semismooth in the sense of Gowda [16]) there. Further, this mapping is almost everywhere strictly differentiable and the semismooth derivative coincides with the generalized Jacobian almost everywhere in $\Omega$. Moreover, its generalized Jacobian is contained in another semismooth derivative obtained from the original one by performing some closure operation and taking the convex hull of the values. For set-valued mappings we introduce a generalization of the semismooth property (denoted $\Gamma$-ss property) and the so-called SCD-ss property. These new notions enable us, among other things, to analyze the semismoothness of solution maps to inclusions appearing in (1.3) by examination of the underlying mapping $F$. Note that, up to now, only results on the semismoothness property of single-valued solution mappings given implicitly by single-valued equations are available, cf. [37, 31, 16, 25, 21].
As already observed by Chen, Sun and Zhang in the very recent paper [5], there are some relationships between G-semismoothness and the semismooth property. We extend these results by showing that a graphically Lipschitzian mapping (i.e., a multifunction whose graph coincides with the graph of a Lipschitz single-valued mapping after a change of coordinates) is SCD-ss if and only if the transformed mapping is G-semismooth. As a byproduct we obtain that an SCD-ss mapping is almost everywhere strictly proto-differentiable and the SC derivative is almost everywhere a singleton showing the easiness of its computation there.
Finally, on the basis of the above theoretical results, we will derive for problem (1.3) two implementable algorithmic schemes, enabling us the computation of a semismooth derivative of $\theta$ via either the limiting coderivative or the SC derivatives of $F$. The respective oracles will then, in cooperation with a suitable bundle method, generate stationary points even in some situations outside the classical area of applicability of ImP. This is illustrated by an academic example. In a subsequent paper we intend to work out this approach to several instances of (1.3) having a practical relevance and to demonstrate its efficiency by numerical experiments.
@@
Let us mention that for single-valued mappings there is a close relation between the semismoothness property and conservative derivatives, the latter being explored in a recent paper by Bolte and Pauwels [2]. In fact, it was shown in [7] that both notions are equivalent for semialgebraic maps. As semismooth derivatives also obey a chain rule like conservative Jacobians, the considerations about automatic differentiation from [2] carry over to our setting. Recently, a conservative Jacobian for solution maps given implicitly by equations was established in [3]. In [4], a conservative Jacobian for the solution map of a parametric monotone inclusion was obtained by rewriting the inclusion as an equivalent fixed-point equation for the forward-backward map by means of the resolvent. Though the resolvents are explicitly known for a lot of monotone mappings, it is not clear how a conservative Jacobian of the resolvent can be computed in the general case. In fact, the resolvent itself may be considered as the solution map of a shifted inclusion. On the contrary, our approach is based on the differentiation of the underlying set-valued mapping by SC derivatives.
@@
The paper is organized in the following way. Section 2 is divided into several subsections with the first two containing the basic notions of variational analysis and generalized differentiation. In Subsection 2.3 we describe a general principle, analogous to the construction of Clarke generalized Jacobian, how to generate from an arbitrary multifunction another multifunction having some favorable properties. In the next subsection we recall the basic properties of SCD (subspace containing derivative) mappings. In Subsection 2.5 we recall the semismooth property of set-valued mappings and its extension to SCD mappings.
@@
The main new results are presented in Sections 3–5. Section 3 is devoted to the semismoothness property of single-valued mappings with respect to certain multifunctions. This concept is commonly used in semismooth Newton methods for solving nonsmooth equations, in particular in infinite dimensions. However, in contrast to these applications, we do not consider the case that semismoothness only holds at the solution, but on the whole domain. There is a close relation with the seminal work of Norkin [29, 30] on generalized differentiation of real-valued functions. In Section 4 we present important relationships between the G-semismooth and the SCD-ss property for graphically Lipschitzian multifunctions. Section 5 concerns the semismooth property of solution maps to inclusions. We establish the semismooth property of general solution maps and the semismooth property for continuous localizations and selections, respectively. Finally it will be shown that the analysis considerably simplifies for SCD mappings.
@@
The following notation is employed. Given a linear subspace $L\subseteq\mathbb{R}^{n}$, $L^{\perp}$ denotes its orthogonal complement. Further, given a multifunction $F$, $\mathrm{gph}\,F:=\{(x,y)\,\big{|}\,y\in F(x)\}$ stands for its graph and $\mathrm{dom}\,F:=\{x\,\big{|}\,F(x)\neq\emptyset\}$ denotes its domain. For an element $u\in\mathbb{R}^{n}$, $\|u\|$ denotes its Euclidean norm and ${\cal B}_{\delta}(u)$ denotes the closed ball around $u$ with radius $\delta$. The closed unit ball is denoted by ${\cal B}$. In a product space we use the norm $\|(u,v)\|:=\sqrt{\|u\|^{2}+\|v\|^{2}}$. The space of all $m\times n$ matrices is denoted by $\mathbb{R}^{m\times n}$. For a matrix $A\in\mathbb{R}^{m\times n}$ , we employ the operator norm $\|A\|$ with respect to the Euclidean norm and we denote the range of $A$ by ${\rm rge\;}A$. Given a set $\Omega\subset\mathbb{R}^{s}$, we define the distance of a point $x$ to $\Omega$ by ${\rm dist}(x,\Omega):=\inf\{\|y-x\|\,\big{|}\,y\in\Omega\}$ and the indicator function is denoted by $\delta_{\Omega}$. Further we define for a mapping $F:\Omega\to\mathbb{R}^{m}$, $\Omega\subset\mathbb{R}^{n}$ open, and $x\in\Omega$ the moduli
@@ 2.1 Nonsmooth analysis
Given a parameterized family $C_{t}$ of subsets of a metric space, where $t$ also belongs to a metric space, we define the upper (outer) and lower(inner) limits in the sense of Painlevé–Kuratowski as
$\mathop{{\rm Lim}\,{\rm sup}}_{t\to\bar{t}}C_{t}:=\{x\,\big{|}\,\liminf_{t\to\bar{t}}\,{\rm dist}(x,C_{t})=0\},\ \liminf_{t\to\bar{t}}C_{t}:=\{x\,\big{|}\,\limsup_{t\to\bar{t}}\,{\rm dist}(x,C_{t})=0\}.$
@@
If $F$ is locally Lipschitz, by Rademacher’s Theorem the set $\Omega\setminus{\cal O}_{F}$ has Lebesgue measure 00. Since the Jacobians $\nabla F(x)$, $x\in{\cal O}_{F}$, are also locally bounded by the Lipschitz constant, we obtain ¯$\overline{\nabla}F(x)\subset\mathrm{conv}\,\overline{\nabla}F(x)$, $x\in\Omega$. Further, we say that $F$ is semismooth at $\bar{x}\in\Omega$ if $F$ is directionally differentiable at $\bar{x}$ and
$\lim_{{\textstyle{x\mathop{\rightarrow}\limits^{\Omega}\bar{x}\atop x\not=\bar{x}}}}\frac{\sup\{\|F(x)-F(\bar{x})-A(x-\bar{x})\|\,\big{|}\,A\in\mathrm{conv}\,\overline{\nabla}F(x)\}}{\|x-\bar{x}\|}=0.$
Note that this definition of semismoothness is different but equivalent to the original ones by Mifflin [26] (for the scalar case) and by Qi and Sun [34]. If the requirement of directional differentiability is dropped we arrive at the semismooth property as defined by Gowda [16], which is also called G-semismoothness in the literature.
We say that $F$ is strictly differentiable at $\bar{x}\in\Omega$ if $F$ is Lipschitz near $\bar{x}$ and there is an $m\times n$ matrix $A$ such that
$\lim_{{\textstyle{x\to\bar{x}\atop t\downarrow 0}}}\frac{F(x+tu)-F(x)}{t}=Au,\ u\in\mathbb{R}^{n}.$
For a locally Lipschitz function $f:\Omega\to\mathbb{R}$, $\Omega\subset\mathbb{R}^{n}$ open, we can define the Clarke subdifferential as
$\partial^{c}f(x):=\mathrm{conv}\,\overline{\nabla}f(x)^{T}:=\{g^{T}\,\big{|}\,g\in\mathrm{conv}\,\overline{\nabla}f(x)\},\ x\in\Omega.$
The elements of $\partial^{c}f(x)$ are called subgradients and are column vectors.
@@
    $\lim_{x{\mathop{\longrightarrow}\limits^{{\cal O}_{F}}}\bar{x}}\nabla F(x)$ exists.
@@
ad (i) $\Rightarrow$ (iii): If $F$ is strictly differentiable at $\bar{x}$ then $\nabla F$ is continuous at $\bar{x}$ relative to ${\cal O}_{F}$, cf. [35, Exercise 9.64]. Thus $\lim_{x{\mathop{\longrightarrow}\limits^{{\cal O}_{F}}}\bar{x}}\nabla F(x)$ exists (and equals to $\nabla F(\bar{x})$).
@@
ad (ii) $\Rightarrow$ (i): Let ¯$\overline{\nabla}F(\bar{x})=\{A\}$. For every $i=1,\ldots,m$, we conclude from [6, Theorem 2.6.6] that the $i$-th component of $F$, $i=1,\ldots m$, satisfies ¯$\bar{\nabla}F_{i}(\bar{x})=\{A_{i}\}$, where $A_{i}$ denotes the $i$-th row of $A$. Thus $\partial^{c}F_{i}(\bar{x})$ is a singleton and therefore $F_{i}$ is strictly differentiable at $\bar{x}$ by [6, Proposition 2.2.4]. Since every component $F_{i}$, $i=1,\ldots m$, is strictly differentiable at $\bar{x}$, so is $F$ as well. ∎